Find the sum of: (a) 137 and – 354 (b) – 52 and 52 (c) – 312, 39 and 192 (d) – 50, – 200 and 300
 
owerbhierhitom,niomnrimmoemhnteioanmeioam,irohomtei

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

(a) 137 and -354 (137) + (-354)
= (317) + (-137) + (-217)
= 0 + (-217) [(317) + (-317) = 0]
= (-217)

(b) -52 and 52
(-52) + (+52) = 0
[(-a) + (+a) = 0]

(c) -312, 39 and 192

(-312) + (+39) + (+192)
= (-231) + (-81) + (+39) + (+192)
= (-231) + (-81) + (+231)
= (-231) + (+231) + (-81)
= 0 + (-81) [(-a) + (+a) = 0]
= -81

(d) -50, -200 and 300
(-50) + (-200) + (+300)
= (-50) + (-200) + (+200) + (+100)
= (-50) + 0 + (+100) [(-a) + (+a) = 0]
= (-50) + (+100)
= (-50) + (+50) + (+50)
= 0 + (+50) [(-a) + (+a) = 0]
=50
Thanks